OpenStudy (anonymous):

A test consists of multiple choice questions worth 2 points and essay questions worth 8 points. The test has 29 questions that are worth a total of 100 points. Use a model to determine how many multiple choice questions are on the test. 16 19 22 25

OpenStudy (anonymous):

Let \(x\) = Number of 2 point questions. \(29-x\) = Number of 8 point questions. \[2x+8(29-x)=100\] \[2x+232-8x=100\] \[6x=132\] \[x=22\] \[29-x=7\] Number of 2 point questions = \(22\) Number of 8 point questions = \(7 \)

OpenStudy (anonymous):

Thanks (:

OpenStudy (anonymous):

You're welcome :-)